OpenStudy (anonymous):

-7(t-9)-(t+4)=4 could someopne help me with this?

jimthompson5910 (jim_thompson5910):

Hint: -7(t-9)-(t+4)=4 -7t+63-t-4=4 -8t+59=4 I'll let you finish

OpenStudy (anonymous):

i dont get how you did that ?

jimthompson5910 (jim_thompson5910):

where are you stuck

OpenStudy (anonymous):

on all of it lol. where does the 63 come from

jimthompson5910 (jim_thompson5910):

Step 2) I distributed Step 3) I combined like terms

jimthompson5910 (jim_thompson5910):

the 63 comes from the fact that -7 times -9 = 63

OpenStudy (anonymous):

i dont get how thats even possible though because wouldnt u hav eto do the t-9 then multiply it by 7?

jimthompson5910 (jim_thompson5910):

-7(t-9) -7 * t + (-7)*(-9) ... distribute -7t + 63 ... multiply

OpenStudy (anonymous):

that makes absolutely no sense to me

jimthompson5910 (jim_thompson5910):

alright, then I suggest you read up on the distributive property http://www.algebrahelp.com/lessons/simplifying/distribution/

OpenStudy (anonymous):

thank u

OpenStudy (mathstudent55):

@mzroze Perhaps this will help you understand the distributive property. 

OpenStudy (mathstudent55):

Similarly, -7(t - 9) = -7(t + -9) = (-7)(t) + (-7)(-9) = -7t + 63 as jim explained

OpenStudy (anonymous):

my milkshake brings all the boys to the yard

OpenStudy (anonymous):

So is the answer t=63.

OpenStudy (mathstudent55):

No, that was just an explanation of the distributive property. -7(t - 9) - (t + 4) = 4 -7t + 63 - t - 4 = 4 -8t + 59 = 4 -8t = -55 t = 55/8